OpenStudy (anonymous):

The speed of a stream is 2 mph. A boat travels 4 miles upstream in the same time it takes to travel 8 miles downstream. What is the speed of the boat in still water?

OpenStudy (anonymous):

sry girl cnt help u

OpenStudy (anonymous):

You sure you got the stream right I mean isn't it 3 mph?

OpenStudy (anonymous):

Ok so first lets try to put this problem into mathematical terms. We know that distance=rate*time, or d=rt, where r is the speed in still water. Also, we know that the stream is moving at 2mph, so if its going upstream the speed will decrease by 2, and if it is going downstram it will increase by 2. So we have: 4=(r-2)t 8=(r+2)t If we solve for t in the second equation, we get t=8/(r+2). Now we can substitute that into the first equation and solve for r: 4=(r-2)*[8/(r+2)] 4=(8r-16)/(r+2) Now we multiply both sides by r+2: 4(r+2)=8r-16) 4r+8=8r-16 4r=24 r=6 Therefore, the speed in still water is 6 mph.

OpenStudy (anonymous):

Thanks!